# A Study On Mixup-inspired Augmentation Methods For Software Vulnerability Detection

### v1

@@ Abstract.
Various Deep Learning (DL) methods have recently been utilized to detect software vulnerabilities. Real-world software vulnerability datasets are rare and hard to acquire as there’s no simple metric for classifying vulnerability. Such datasets are heavily imbalanced, and none of the current datasets are considered huge for DL models. To tackle these problems a recent work has tried to augment the dataset using the source code and generate realistic single-statement vulnerabilities which is not quite practical and requires manual checking of the generated vulnerabilities. In this regard, we aim to explore the augmentation of vulnerabilities at the representation level to help current models learn better which has never been done before to the best of our knowledge. We implement and evaluate the 5 augmentation techniques that augment the embedding of the data and recently have been used for code search which is a completely different software engineering task. We also introduced a conditioned version of those augmentation methods, which ensures the augmentation does not change the vulnerable section of the vector representation. We show that such augmentation methods can be helpful and increase the f1-score by up to 9.67%, yet they cannot beat Random Oversampling when balancing datasets which increases the f1-score by 10.82%!
@@ 1. INTRODUCTION
In Software Engineering and Development, software vulnerabilities are widespread which are the cause of many security risks in the industry. To help with this problem, researchers have tried to come up with models that can predict and classify the vulnerability of code pieces (Scandariato et al., 2014; Hin et al., 2022; Li et al., 2021; Wang et al., 2023; Fu and Tantithamthavorn, 2022; Li et al., 2018; Zhou et al., 2019; Chakraborty et al., 2022). However, all of these works face the reality of being unable to perform well in real-life scenarios due to a shortage of data. Chakraborty et al.(Chakraborty et al., 2022) revealed a substantial 73% average drop in F1 scores for state-of-the-art deep learning-based vulnerability detection (DLVD) models when dealing with real-world imbalanced data. This performance decline poses a challenge to the practical application of DLVD, prompting the exploration of strategies like data sampling (Yang et al., 2023), data augmentation and generation (Nong et al., 2023, 2024; Daneshvar et al., 2024), cost-sensitive learning (Nguyen et al., 2020), and ensemble methods (Seiffert et al., 2009).
@@
To deal with the shortage of vulnerable samples, Nong, et al. (Nong et al., 2023, 2024) proposed limited solutions (i.e. VGX and Vulgen) for generating single statement vulnerabilities using a deep learning model and a data mining-based method. Yang et al. (Yang et al., 2023) explored the effect of different sampling methods both at the raw level and latent level, and found that the rudimentary random oversampling (ROS) achieves the best performance and more complex methods such as SMOTE  (Chawla et al., 2002) cannot keep up with ROS. Recently, Daneshvar et al. (Daneshvar et al., 2024) showed that all the previous methods have limited effectiveness and similar to SMOTE, Vulgen and VGX cannot beat ROS. They also proposed a vulnerable code generation pipeline, heavily based on prompting LLMs, that could beat ROS but was costlier than the previous methods, and does not work without using powerful LLMs.
@@
Representation-level augmentation methods’ effectiveness has been shown for various Software Engineering tasks, namely, Code Search (Li et al., 2022), Problem Classification (Dong et al., 2023a, b), Bug Detection (Dong et al., 2023a, b), Authorship Attribution (Dong et al., 2023b), and Clone Detection (Dong et al., 2023b). However, the efficacy of such methods remains unexplored for the task of vulnerability detection. While this task is not directly related to Bug Detection, it is closer than others, yet significantly more challenging. Vulnerable code functioning correctly adds complexity, requiring increased expertise and focus in real-life scenarios. Additionally, vulnerability detection datasets are smaller and more sparse. It is important to note that the nature of vulnerability datasets and the complexity of the task itself, make it imprudent to assume the effectiveness of these methods for vulnerability detection without conducting systematic evaluation.
@@
Hence, following the recent work on augmenting data for more common code tasks by Dong et al. (Dong et al., 2023a) and on code search by Li et al. (Li et al., 2022), we explore and test if the text-based embedding augmentation techniques, which are all influenced by the idea of MixUp (Guo et al., 2019), help with the training of token-based vulnerability detection models, such as LineVul (Fu and Tantithamthavorn, 2022), and whether they can beat ROS. We also use VGX as another baseline and show that these methods are more suitable than VGX for large-scale data augmentation.
@@
In this paper, we explore the effect of five representation-level augmentation methods (Linear Interpolation, Stochastic Perturbation, Linear Extrapolation, Binary Interpolation, and Gaussian Scaling). Uniquely adapting these techniques for vulnerability detection, we exclusively augment the vulnerable items. Additionally, we explore a variant of these methods by conditioning the augmentation, keeping the vulnerable part of the code fixed. This approach is non-intrusive, as it does not necessitate extensive source code analysis. We evaluate the data augmentation methods on LineVul (Fu and Tantithamthavorn, 2022), a SOTA token-based DLVD, with BigVul (Fan et al., 2020) dataset, and compare the results. The results show that the conditioned variation performs better in general as it provides a more targeted enhancement to vulnerability detection by ensuring the generated embeddings contain vulnerable lines. However, ROS generates a better performance. More specifically, ROS boosts the f1-score performance by 10.82% while the best-performing augmentation method (i.e. naive stochastic perturbation) boosts the performance by 9.96%.
In summary, the contributions of our paper include:
@@
  • To our knowledge, we conducted the first systematic and thorough study to understand how representation-level augmentation affects DLVD. We examined the impact of five data augmentation methods on a leading token-based DLVD approach, utilizing one of the largest datasets available that provide line-level information. Line-level information is crucial, particularly for the conditioned approach.
@@
  • Our code is available here for replication and further improvements.
2. Background and Related Work
This section introduces the background and related work to DLVD and representation-level augmentation.
@@ 2.1. Overview of deep learning-based vulnerability detection
Generally, DLVD is done in three phases: feature extraction, model training, and model deployment. In the feature extraction phase, various features are extracted from the code that can effectively capture the semantic and syntactic properties of the code. Following this, the extracted features are embedded into a real-valued vector, representing the code snippet. Subsequently, these representations are used to train the model that predicts the vulnerability of a representation. For instance, LineVul (Fu and Tantithamthavorn, 2022) is a state-of-the-art method that leverages CodeBERT (Feng et al., 2020) for embedding code snippets (both vulnerable and clean samples). Finally, in the deployment phase, the same feature extraction and embedding techniques are used to transform code snippets of software systems and feed that to the model to predict the vulnerability.
@@ 2.2. Token-based Vulnerability Detection and LineVul
Derived from the extraction and embedding processes, two classes of DLVD can be identified: Token-based and Graph-based, in our study, we focus on the former. In token-based DLVD approaches, each code snippet is transformed into a sequence of learnable vectors known as tokens. The token representations of the code are acquired through text embedding techniques such as Word2Vec (Mikolov et al., 2013) and BERT  (Devlin et al., 2019). For instance, LineVul (Fu and Tantithamthavorn, 2022) is a recent transformer-based (Vaswani et al., 2017) state-of-the-art method that predicts the vulnerability of functions and can locate the vulnerable lines.
@@ 2.3. Representation-level Augmentation
Data augmentation for source code can be approached in various ways  (Dong et al., 2023b; Bui et al., 2021; Dong et al., 2023a). However, many of these methods are designed specifically for source code and necessitate source code analysis. This process is time-consuming and involves altering the code meticulously to ensure the preservation of code semantics, such as Variable Renaming and Switch to If (Bui et al., 2021). An alternative to augmenting source code directly is representation-level augmentation. This method involves augmenting data after the embedding process, and manipulating the representation vectors using mathematical processes to generate new data points. Representation-level augmentation can be applied to both text and graph data.
In this section, we delve into the five token-based representation-level augmentation methods following previous studies (Li et al., 2022; Dong et al., 2023b, a). These methods encompass the entire spectrum of representation-level augmentation explored for source code at the text level. All 5 of those methods can be expressed using the following general format:
(1) $h^{+}=\alpha\odot h+\beta\odot h^{\prime}$
where $h^{+}$ is the augmented data, $\alpha$ and $\beta$ are coefficients determined by the method, $h$ and $h^{\prime}$ are two data points in their representation vector format, drawn from the original dataset. Based on the general format these augmentation methods are described as follows:
  • Linear Interpolation (LI) This is the initial method used for representation-level augmentation of source code and was inspired by SenMixup (Guo et al., 2019) in NLP, hence it may be called Mixup (Zhang et al., 2017) for source code as well. In this method, $\alpha$ is sampled from a uniform distribution $\alpha\sim U(a,1.0)$ for large values of $a$ and $\beta=1-\alpha$, and $h\neq h^{\prime}$. The application of this method is depicted in Figure 1, and our conditioned variation of it, in Figure 2.
  • Stochastic Perturbation (SP) This method randomly deactivates features in the representation vector and usually is implemented using DropOut, which drops features with a probability of p and then multiplies the resulting vector by a factor of $\frac{1}{(1-p)}$. Hence, in the case of Stochastic Perturbation, $\alpha$ is sampled from a Bernoulli distribution $B(p)$, and $p$ is the probability parameter. Consequently, $\alpha\in\{0,\frac{1}{1-p}\}$, $\beta=\frac{1}{1-p}-\alpha$ and $h^{\prime}=0$.
  • Linear Extrapolation (LE) This method resembles Linear Interpolation, but it generates data points outside the hypersphere instead of inside it. In other words, the only difference is $\alpha\sim U(1.0,a)$, for small values of $a$, so that the generated data point is close to one of the original data points used for augmentation.
  • Binary Interpolation (BI) This method involves randomly exchanging certain features with those from another selected sample; therefore, It is similar to Stochastic Perturbation but instead of deactivating features, it replaces them with features from another sample. Hence, in the case of Binary Interpolation $\alpha$ is sampled from a Bernoulli distribution $B(p)$, $\alpha\in\{0,1\}^{e}$, $\beta=1-\alpha$, and $h\neq h^{\prime}$.
  • Gaussian Scaling (GS) This method produces scaling coefficients for each feature in the representation vector, akin to a form of perturbation noise. In the case of this method, $h=h^{\prime}$, $\alpha=1$, and $\beta\sim N(1,\sigma)$ with small values of $\alpha$.
@@
3. Experiment Design
@@ 3.1. Blind augmentation
We refer to applying the 5 augmentation methods, introduced in section 2.3, to vulnerable samples of the BigVul dataset as Blind Augmentation. This is to emphasize that the vulnerable section in each vulnerable section may be overwritten or removed.
@@ 3.2. Conditioned augmentation
Aside from blindly applying 5 augmentation approaches, we also investigate the effect of fixing the vulnerable sections (i.e. conditioned augmentation) while using such augmentation methods. This involved intelligently excluding the tokens in the embedding corresponding to the specified vulnerable statement(s) within the dataset from the augmentation process. Utilizing Byte-Pair Encoding (BPE) for transforming code into tokens, we could easily identify the transformed vulnerable line in the embedding. Despite minor variations in the initial and final tokens due to the nature of BPE, we systematically searched for the sequence of tokens in the code embedding. In case of failure, we recursively excluded the first and last tokens in the line embedding until success. Subsequently, we replaced the corresponding tokens of the vulnerable line(s) to ensure the final generated vector encapsulated the vulnerability.
@@ 3.3. Implementation Details
To evaluate different data augmentation methods, we used the Bigvul (Fan et al., 2020) dataset, shown in Table 1, for training and testing Linevul (Fu and Tantithamthavorn, 2022). Specifically, we used the original training set of Bigvul and augmented the vulnerable samples of the training set that contained the vulnerable line information so that it would be a fair comparison between the naive and conditioned augmentation methods. We trained Linevul with the optimal hyperparameters mentioned in Table 2, with embedding block size and feature embedding size similar to the original (Fu and Tantithamthavorn, 2022). Feature embedding size corresponds to the vectors’ size at the end of RoBERTa (Liu et al., 2019) before being fed to the final classifier. We also used the validation set to select the best checkpoint during training for testing. To compare these methods we employed the four common evaluation metrics (Powers, 2011), namely recall, precision, F1-score, and Area Under The Curve (AUC), in line with previous studies (Yang et al., 2023; Li et al., 2018; Wang et al., 2023; Rajbahadur et al., 2022, 2021; Daneshvar et al., 2024).
@@
Batch Size #Epochs Learning Rate Block Size Feature Embedding Size
@@
Aside from the two baselines of ROS and VGX, we investigate the application of the 5 methods outlined in Section 2.3 in both settings (i.e. naive and conditioned), the details of which are provided below:
  • Linear Interpolation (LI) We employed this method with elements of $\alpha$ sampled from the uniform distribution $\alpha\sim U(0.9,1.0)$. Figure 1 illustrates an example of this method with the sampled $\alpha$ set to 0.95.
  • Stochastic Perturbation (SP) In alignment with prior studies, we utilized DropOut for the implementation of this method with a $p=0.1$.
  • Linear Extrapolation (LE) We employed this method with elements of $\alpha$ sampled from the uniform distribution $\alpha\sim U(1.0,1.1)$.
  • Binary Interpolation (BI) For the implementation of this method, we opted to alter only 25% of the cells each time, replacing them with those from another vulnerable sample, following the approach outlined by Li et al.(Li et al., 2022).
  • Gaussian Scaling (GS) In the case of this method, we sampled $\beta$ from a Gaussian distribution $\beta\sim N(1,0.1)$.
  • Random Oversampling (ROS) We used ROS as a baseline as it can benefit DLVD and is not easy to beat (Yang et al., 2023; Daneshvar et al., 2024). Hence, we used ROS to sample the raw vulnerable samples randomly to get a balanced dataset.
  • VGX Similar to ROS, we used VGX (Nong et al., 2024) as another baseline and sampled the same number of vulnerable samples (i.e. 129690 items) from the 896K items provided by the original paper (Nong et al., 2024).
For all of the mentioned augmentation methods in both settings, vulnerable samples were used to generate new ones; therefore, all the generated items were labeled as vulnerable. We augmented the data at a rate of 23 using all the aforementioned methods. This implies that the augmented dataset comprises the original items along with augmented data, amounting to 22 times the number of vulnerable samples. This approach was adopted to achieve a balanced dataset.

### v2

@@ Abstract.
Various deep learning (DL) methods have recently been utilized to detect software vulnerabilities. Real-world software vulnerability datasets are rare and hard to acquire, as there is no simple metric for classifying vulnerability. Such datasets are heavily imbalanced, and none of the current datasets are considered huge for DL models. To tackle these problems, a recent work has tried to augment the dataset using the source code and generate realistic single-statement vulnerabilities, which is not quite practical and requires manual checking of the generated vulnerabilities. In this paper, we aim to explore the augmentation of vulnerabilities at the representation level to help current models learn better, which has never been done before to the best of our knowledge. We implement and evaluate five augmentation techniques that augment the embedding of the data and have recently been used for code search, which is a completely different software engineering task. We also introduced a conditioned version of those augmentation methods, which ensures the augmentation does not change the vulnerable section of the vector representation. We show that such augmentation methods can be helpful and increase the F1-score by up to 9.67%, yet they cannot beat Random Oversampling when balancing datasets, which increases the F1-score by 10.82%.
@@ 1. Introduction
In software engineering and development, software vulnerabilities are widespread, which are the cause of many security risks in the industry. To help with this problem, researchers have tried to come up with models that can predict and classify the vulnerability of code pieces (Scandariato et al., 2014; Hin et al., 2022; Li et al., 2021; Wang et al., 2023; Fu and Tantithamthavorn, 2022; Li et al., 2018; Zhou et al., 2019; Chakraborty et al., 2022). However, all of these works face the reality of being unable to perform well in real-life scenarios due to a shortage of data. Chakraborty et al. (Chakraborty et al., 2022) revealed a substantial 73% average drop in F1-scores for state-of-the-art (SOTA) deep learning-based vulnerability detection (DLVD) models when dealing with real-world imbalanced data. This performance decline poses a challenge to the practical application of DLVD, prompting the exploration of strategies like data sampling (Yang et al., 2023), data augmentation and generation (Nong et al., 2023, 2024; Daneshvar et al., 2024), cost-sensitive learning (Nguyen et al., 2020), and ensemble methods (Seiffert et al., 2009).
@@
To deal with the shortage of vulnerable samples, Nong et al. (Nong et al., 2023, 2024) proposed limited solutions (i.e., VGX and Vulgen) for generating single statement vulnerabilities using a deep learning model and a data mining-based method. Yang et al. (Yang et al., 2023) explored the effect of different sampling methods both at the raw level and latent level, and found that the rudimentary Random Oversampling (ROS) achieves the best performance and more complex methods, such as SMOTE  (Chawla et al., 2002) cannot keep up with ROS. Recently, Daneshvar et al. (Daneshvar et al., 2024) showed that all the previous methods have limited effectiveness and similar to SMOTE, Vulgen and VGX cannot beat ROS. They also proposed a vulnerable code generation pipeline, heavily based on prompting LLMs, that could beat ROS but was costlier than the previous methods, and does not work without using powerful LLMs.
@@
Representation-level augmentation methods’ effectiveness has been shown for various software engineering tasks—namely, code search (Li et al., 2022), problem classification (Dong et al., 2023a, b), bug detection (Dong et al., 2023a, b), authorship attribution (Dong et al., 2023b), and clone detection (Dong et al., 2023b). However, the efficacy of such methods remains unexplored for the task of vulnerability detection. While this task is not directly related to bug detection, it is closer than others, yet significantly more challenging. Vulnerable code functioning correctly adds complexity, requiring increased expertise and focus in real-life scenarios. Additionally, vulnerability detection datasets are smaller and more sparse. It is important to note that the nature of vulnerability datasets and the complexity of the task itself makes it imprudent to assume the effectiveness of these methods for vulnerability detection without conducting a systematic evaluation.
@@
Hence, following the recent work on augmenting data for more common code tasks by Dong et al. (Dong et al., 2023a) and on code search by Li et al. (Li et al., 2022), we explore and test if the text-based embedding augmentation techniques. These techniques are all influenced by the idea of MixUp (Guo et al., 2019), help with the training of token-based vulnerability detection models—such as LineVul (Fu and Tantithamthavorn, 2022)—and whether they can beat ROS. We also use VGX as another baseline and show that these methods are more suitable than VGX for large-scale data augmentation.
@@
In this paper, we explore the effect of five representation-level augmentation methods (Linear Interpolation, Stochastic Perturbation, Linear Extrapolation, Binary Interpolation, and Gaussian Scaling). Uniquely adapting these techniques for vulnerability detection, we exclusively augment the vulnerable items. Additionally, we explore a variant of these methods by conditioning the augmentation, keeping the vulnerable part of the code fixed. This approach is non-intrusive as it does not necessitate extensive source code analysis. We evaluate the data augmentation methods on LineVul (Fu and Tantithamthavorn, 2022) (a SOTA token-based DLVD) with BigVul (Fan et al., 2020) dataset, and compare the results. The results show that the conditioned variation performs better in general, as it provides a more targeted enhancement to vulnerability detection by ensuring the generated embeddings contain vulnerable lines. However, ROS generates a better performance. More specifically, ROS boosts the F1-score performance by 10.82% while the best-performing augmentation method (i.e., naive Stochastic Perturbation) boosts the performance by 9.96%. Hence, key contributions of our paper include:
@@
  • To our knowledge, we conducted the first systematic and thorough study to understand how representation-level augmentation affects DLVD. We examined the impact of five data augmentation methods on a leading token-based DLVD approach, utilizing one of the largest datasets available that provides line-level information. Line-level information is crucial, particularly for the conditioned approach.
@@
  • We comply with the SIGSOFT Open Science Policy by putting our code in a permanent repository—namely, Zenodo—to the extent ethically and practically possible for replication and further improvements.
The remainder of this paper is organized as follows. The next section introduces background and related works to DLVD and representation-level augmentation. Section 3 describes our experimental design, and Section 4 shows results. Sections 5, 6 and 7 provide discussions, conclusions, and code availability, respectively.
2. Background and Related Works
@@ 2.1. Overview of Deep Learning-Based Vulnerability Detection
Generally, DLVD is done in three phases: feature extraction, model training, and model deployment. In the feature extraction phase, various features are extracted from the code that can effectively capture the semantic and syntactic properties of the code. Following this, the extracted features are embedded into a real-valued vector, representing the code snippet. Subsequently, these representations are used to train the model that predicts the vulnerability of a representation. For instance, LineVul (Fu and Tantithamthavorn, 2022) is a SOTA method that leverages CodeBERT (Feng et al., 2020) for embedding code snippets (both vulnerable and clean samples). Finally, in the deployment phase, the same feature extraction and embedding techniques are used to transform code snippets of software systems and feed them to the model to predict the vulnerability.
@@ 2.2. Token-Based Vulnerability Detection and LineVul
Derived from the extraction and embedding processes, two classes of DLVD can be identified: Token-based and graph-based. In our study, we focus on the former. In token-based DLVD approaches, each code snippet is transformed into a sequence of learnable vectors known as tokens. The token representations of the code are acquired through text embedding techniques such as Word2Vec (Mikolov et al., 2013) and BERT (Devlin et al., 2019). For instance, LineVul (Fu and Tantithamthavorn, 2022) is a recent transformer-based (Vaswani et al., 2017) SOTA method that predicts the vulnerability of functions and can locate the vulnerable lines.
@@ 2.3. Representation-Level Augmentation
Data augmentation for source code can be approached in various ways  (Dong et al., 2023b; Bui et al., 2021; Dong et al., 2023a). However, many of these methods are designed specifically for source code and necessitate source code analysis. This process is time-consuming and involves altering the code meticulously to ensure the preservation of code semantics, such as Variable Renaming and Switch to If (Bui et al., 2021). An alternative to augmenting source code directly is representation-level augmentation. This method involves augmenting data after the embedding process and manipulating the representation vectors using mathematical processes to generate new data points. Representation-level augmentation can be applied to both text and graph data.
@@
In this section, we delve into the five token-based representation-level augmentation methods following previous studies (Li et al., 2022; Dong et al., 2023b, a). These methods encompass the entire spectrum of representation-level augmentation explored for source code at the text level. All five methods can be expressed using the following general format:
(1) $h^{+}=\alpha\odot h+\beta\odot h^{\prime}$
where $h^{+}$ is the augmented data, $\alpha$ and $\beta$ are coefficients determined by the method, $h$ and $h^{\prime}$ are two data points in their representation vector format, drawn from the original dataset. Based on the general format, these augmentation methods are described as follows:
  • Linear Interpolation (LI), which is the initial method used for representation-level augmentation of source code and was inspired by SenMixup (Guo et al., 2019) in NLP. Hence, it may be called Mixup (Zhang et al., 2017) for source code as well. In this method, $\alpha$ is sampled from a uniform distribution $\alpha\sim U(a,1.0)$ for large values of $a$ and $\beta=1-\alpha$, and $h\neq h^{\prime}$. The application of this method is depicted in Figure 1, and our conditioned variation of it in Figure 2.
  • Stochastic Perturbation (SP), which randomly deactivates features in the representation vector. It is usually implemented using DropOut, which drops features with a probability of $p$ and then multiplies the resulting vector by a factor of $\frac{1}{1-p}$. Hence, in the case of SP, $\alpha$ is sampled from a Bernoulli distribution $B(p)$, and $p$ is the probability parameter. Consequently, $\alpha\in\left\{0,\frac{1}{1-p}\right\}$, $\beta=\frac{1}{1-p}-\alpha$, and $h^{\prime}=0$.
  • Linear Extrapolation, which resembles LI, but it generates data points outside the hypersphere instead of inside it. In other words, the only difference is $\alpha\sim U(1.0,a)$, for small values of $a$, so that the generated data point is close to one of the original data points used for augmentation.
  • Binary Interpolation, which involves randomly exchanging certain features with those from another selected sample. Hence, it is similar to SP but instead of deactivating features, it replaces them with features from another sample. Hence, in the case of BI, $\alpha$ is sampled from a Bernoulli distribution $B(p)$, $\alpha\in\{0,1\}^{e}$, $\beta=1-\alpha$, and $h\neq h^{\prime}$.
  • Gaussian Scaling, which produces scaling coefficients for each feature in the representation vector, akin to a form of perturbation noise. In this method, $h=h^{\prime}$, $\alpha=1$, and $\beta\sim N(1,\sigma)$ with small values of $\alpha$.
3. Experimental Design
@@ 3.1. Blind Augmentation
We applied the five augmentation methods, as introduced in Section 2.3, to vulnerable samples of the BigVul dataset as Blind Augmentation. This is to emphasize that the vulnerable section in each vulnerable section may be overwritten or removed.
@@ 3.2. Conditioned Augmentation
Aside from blindly applying five augmentation approaches, we also investigated the effect of fixing the vulnerable sections (i.e., conditioned augmentation) while using such augmentation methods. This involved intelligently excluding the tokens in the embedding corresponding to the specified vulnerable statement(s) within the dataset from the augmentation process. Utilizing Byte-Pair Encoding (BPE) for transforming code into tokens, we could easily identify the transformed vulnerable line in the embedding. Despite minor variations in the initial and final tokens due to the nature of BPE, we systematically searched for the sequence of tokens in the code embedding. In case of failure, we recursively excluded the first and last tokens in the line embedding until success. Subsequently, we replaced the corresponding tokens of the vulnerable line(s) to ensure the final generated vector encapsulated the vulnerability.
@@ 3.3. Implementation Details
To evaluate different data augmentation methods, we used the Bigvul (Fan et al., 2020) dataset—as shown in Table 1—for training and testing Linevul (Fu and Tantithamthavorn, 2022). Specifically, we used the original training set of Bigvul and augmented the vulnerable samples of the training set that contained the vulnerable line information so that it would be a fair comparison between the naive and conditioned augmentation methods. We trained Linevul with the optimal hyperparameters mentioned in Table 2, with embedding block size and feature embedding size similar to the original (Fu and Tantithamthavorn, 2022). Feature embedding size corresponds to the vectors’ size at the end of RoBERTa (Liu et al., 2019) before being fed to the final classifier. We also used the validation set to select the best checkpoint during training for testing. To compare these methods, we employed the four common evaluation metrics (Powers, 2011)—namely recall, precision, F1-score, and area under the curve (AUC)—in line with previous studies (Yang et al., 2023; Li et al., 2018; Wang et al., 2023; Rajbahadur et al., 2022, 2021; Daneshvar et al., 2024).
@@
Batch #epochs Learning Block Feature embedding
size rate size size
@@
Aside from the two baselines of ROS and VGX, we investigate the application of the five methods outlined in Section 2.3 in both settings (i.e., naive and conditioned):
  • Linear Interpolation. We employed this method with elements of $\alpha$ sampled from the uniform distribution $\alpha\sim U(0.9,1.0)$. Figure 1 illustrates an example of this method with the sampled $\alpha$ set to 0.95.
  • Stochastic Perturbation. In alignment with prior studies, we utilized DropOut for the implementation of this method with a $p=0.1$.
  • Linear Extrapolation. We employed this method with elements of $\alpha$ sampled from the uniform distribution $\alpha\sim U(1.0,1.1)$.
  • Binary Interpolation. For this method, we opted to alter only 25% of the cells each time, replacing them with those from another vulnerable sample, following the approach outlined by Li et al.(Li et al., 2022).
  • Gaussian Scaling. For this method, we sampled $\beta$ from a Gaussian distribution $\beta\sim N(1,0.1)$.
  • Random Oversampling (ROS). We used ROS as a baseline as it can benefit DLVD and is not easy to beat (Yang et al., 2023; Daneshvar et al., 2024). Hence, we used ROS to sample the raw vulnerable samples randomly to get a balanced dataset.
  • VGX. Similar to ROS, we used VGX (Nong et al., 2024) as another baseline and sampled the same number of vulnerable samples (i.e., 129690 items) from the 896K items provided by the original paper (Nong et al., 2024).
For all of the mentioned augmentation methods in both settings, vulnerable samples were used to generate new ones; therefore, all the generated items were labeled as vulnerable. We augmented the data at a rate of 23 using all the aforementioned methods. This implies that the augmented dataset comprises the original items along with augmented data, amounting to 22$\times$ the number of vulnerable samples. This approach was adopted to achieve a balanced dataset.